Waterways Corporation uses very stringent standard costs in evaluating its manufacturing efficiency. These standards are not “ideal” at this point, but the management is working toward that as a goal. At present, the company uses the following standards.

Materials
Item Per unit Cost
Metal 1 lb. 63¢ per lb.
Plastic 12 oz. $1.00 per lb.
Rubber 4 oz. 88¢ per lb.
Direct labor
Item Per unit Cost
Labor 15 min. $9.00 per hr.
Predetermined overhead rate based on direct labor hours = $4.24


The January figures for purchasing, production, and labor are:

The company purchased 228,200 pounds of raw materials in January at a cost of 79¢ a pound.
Production used 228,200 pounds of raw materials to make 115,000 units in January.
Direct labor spent 18 minutes on each product at a cost of $8.80 per hour.
Overhead costs for January totaled $54,597 variable and $72,000 fixed.

Answer the following questions about standard costs. And list if they are favorable, unfavorable, or neither

What is the materials price variance? Materials price variance $

What is the materials quantity variance? Materials quantity variance $

What is the total materials variance? Total materials variance $

What is the labor price variance? Labor price variance $

What is the labor quantity variance?Labor quantity variance $

What is the total labor variance?Total labor variance $

What is the total overhead variance? Total overhead variance $
Answer #1

ANSWER

1. Materials price variance:

Standard price per unit ($) = (1 * 0.63) + (0.75 * 1) + (0.25 * 0.88) = 0.63 + 0.75 + 0.22 = 1.6 per unit

Material Price variance = Actual quantity * (Standard price - Actual price)

Material price variance = 228,200 * $(0.79 - 1.6) = $184,842(favorable)


2. Materials quantity variance:

Material quantity variance = Standard price * (Actual quantity - Standard quantity)

= $1.6 x (228,200 - 230,000) = $1.6 x 1,800= $2,880(Favorable)


3. Total materials variance:

Total material variance = Price variance + Quantity variance

Total material variance = $184,842(favorable) + $2,880(Favorable) = $187,722(favorable)


4. Labor price variance:

Labor price variance = Actual hours * (Actual rate - Standard rate)

Labor price variance = (115,000 * 18 / 60) * $(8.80 - 9) = 34,500 * $0.2 = $6,900(Favorable)


5. Labor quantity variance:

Labor quantity variance = Standard rate x (Actual hours - Standard hours)

Labor quantity variance = $9 * [(115,000 * 18 / 60) - (115,000 * 15 / 60)] = $51,750 (Unfavorable)


6. Total labor variance:

Total labor variance = Price variance + Quantity variance

Total labor variance = $6,900 (Favorable) + $51,750 (Unfavorable) = $44,850(Unfavorable)

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************